11. Calculate the future value of $7,000 in

a. Four years at an interest rate of 6% per year.

b.Eight years at an interest rate of 6% per year.

C. Four years at an interest rate of 12% per year.

d. Why is the amount of interest earned in part ​(a​) less than half the amount of interest earned in part ​(b​)?
Answer #1

Present value = Future value/(1+i)^n

i = interest rate per period

n= number of periods

a)

Future value = 7000 * (1+6%)^4

= 8837.34

b)

Future value = 7000 * (1+6%)^8

= 11156.94

c)

Future value = 7000 * (1+12%)^4

= 11014.64

d)

Due to the principal of compounding, As money starts to accumulate, interest going to increase